A telephone installer has 115 feet of wire which is to be cut into three pieces. The second piece is 5 feet longer than the first piece, and the length of the third piece is three times the length of the second. Find the length of each piece.
Answer by scott8148(6628)

let x=first

"The second piece is 5 feet longer than the first piece" ___ x+5=second

"the length of the third piece is three times the length of the second" ___ 3(x+5)=third

x+x+5+3(x+5)=115 ___ x+x+5+3x+15=115 ___ 5x+20=115